Trying to help out young friend, I have been out of school for 25yrs and not sure if I am on the right path, young man is going to fail math and really needs help:
If y=x-2, find x in 4x+3y+78
This is what I got but not sure:
4x+3y=78
4(y-3)+3y=78
4y-6+3y=78
7y-6+6=78+6
7y=84
y=12
am I way of? Can't figure out how to get x+?
Please help, Thank you.

Answer by checkley77(12569)

YOU MAKE A MISTAKE BY REPLACING X BY (Y-3)
if y=x-2, find x in 4x+3y=78 IF THIS IS THE EQUATION??? BY REPLAXCING Y BY (X-2)4x+3(x-2)=78
4X+3X-6=78
7X=78+6
7X=84
X=12 ANSWER.